Solve for x
3/4(x + 2) - 3 = 5/6(2x)

Answer by aswathytony(47)


3/4 (x +2) -3 = 5/6 *2x

3/4x +3/4*2 -3 =5/3x
3/4 x -5/3x + 3/2 -3 =0
x ( 3/4 -5/3) +3/2-3 =0
x(9-20/12) + 3-6/3 =0
-11/12 x + -3/3 =0
-11/12 x +-1 =0
-11/12 x =1
x = -12/11